A straight line through the origin meets the parallel lines 4x+2y=9 and 2x+y+6=0 at P and Q respectively. Then the point O divides PQ in the ratio???

Grade:12

2 Answers

Chetan Mandayam Nayakar
12 years ago

Dear Pranavi,

y=kx,(2+k)x=-6,

x=-6/(k+2),y=-6k(k+2),this is Q

now P:(4+2k)x=9

x=9/(4+2k),y=9k/(4+2k)

OP=(9/(4+2k))√(1+k2)

OQ=-(6/(2+k))√(1+k2)

lOP/OQl=l-(9/6)*(k+2)/(2k+4)l=3/4

 

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Pranavi !!!

 


Regards,

Askiitians Experts

CHETAN MANDAYAM NAYAKAR

anahita vn
12 years ago

PO : OQ = 3:4

 2y = -4x + 9

2y = -4x - 12


The y-intercepts of these two lines are +9/2 and -12/2, respectively. The origin divides the segment as

9:12 = 3:4